What must you do to make a 10 mM solution of NaOH when you are given a stock of NaOH that is 1 MM ?
1
Expert's answer
2019-08-14T05:56:56-0400

1 MM solution = 10^6 M solution

1 mM solution = 10^(-3) M solution

The required dilution is 10^6 / 10^(-3) = 10^9

a series of 10-fold dilutions should be made.

1 ml of the initial solution should be transferred to 9 ml of water and mixed. Then, 1 ml of the obtained mixture should be transferred to 9 ml of water. The procedure should be repeated nine times.